OpenStudy (anonymous):

Find the center and vertices of the ellipse. x2 + 4y2 - 4x + 8y + 4 = 0

OpenStudy (anonymous):

x^2 +4y^2-4x+8y+4=0 to clarify

OpenStudy (anonymous):

Complete the squares \[x ^{2}-4x+4+4y ^{2}+8y+4=4\] \[(x-2)^{2}+(2y+2)^{2}=2^{2}\] so the center is (2,-1) If x=2, y=-2 or y=0 If y=-1, x=4 or x=0 (2,-2),(2,0),(4,-1),(0.-1)

OpenStudy (anonymous):

thank you!! :)